# Improving Multimodal Large Language Models Using Continual Learning

### v1

@@ B.3 Evaluation Dataset Preparation
Corresponding to the datasets used to train LLaVA, we evaluate the model on all corresponding datasets. We use the slim [^1] https://github.com/TRI-ML/vlm-evaluation versions of all VQAv2, GQA, Text-VQA and RefCOCO datasets for evaluation, as provided within Prismatic-VLMs [31]. All slim versions of the evaluation sets contain $1024$ examples each, and we use the provided index splits for testing.
B.4 Measuring Performance
To assess performance, we compute the forgetting metric $\Delta$, where a positive value indicates forgetting and a negative value indicates learning the next task enhances the performance of previously acquired tasks (or a backward transfer [40]). The performance change, $\Delta$, on task $t$ after training on task $k$ is defined as:
$\Delta_{t}(k)=\omega_{t}(1)-\omega_{t}(k),\quad\forall t<T$ (1)
where $T$ is the total number of tasks, and $\omega_{t}$ is the harmonic mean of the accuracy values of the evaluation datasets for task $t$. We use the harmonic mean because it is dominated by the accuracy of the worst-performing dataset for the task, thereby emphasizing the need to perform well and avoid forgetting for all of the datasets.
Appendix C Additional Experiments
C.1 LLaVA Model Scaling
Below, we report the VL performance results with varying model scales after LLaVA training. We also again show the linguistic forgetting here for reference. In Figure 4(a), we see that the VL performance of the Soft Targets approach is competitive with the Naive FT approach, especially at the 2.8B scale. We note that the gap in VL performance reduces with the model scale in the LLaVA, and is nearly on par with Naive FT at the highest scales. This is while maintaining a zero to negative loss in linguistic abilities compared to the Naive FT approach (Figure  4(b)).
(a) Avg. VL performance with model scale
(b) Linguistic forgetting with model scale
Figure 4: Vision-language and linguistic forgetting for varying model sizes. We evaluate the average VL performance and linguistic forgetting for Pythia models in the 160M to 2.8B scale, after training on LLaVA. Negative forgetting refers to a positive backward transfer in NL performance after multimodal training.
C.2 Continual LLaVA Model Scaling
Model scaling results are given in Fig. 5. We compare mSGM with Rehearsal, the best method identified for Pythia 410M with naive fine-tuning. Naive fine-tuning leads to a sharp and consistent drop in NLU/NLG performance as continual multi-modal training proceeds. In contrast, mSGM with Rehearsal has little to no NLU/NLG forgetting across all tasks and multiple model scales. For Pythia 1B, mSGM with Rehearsal achieves positive backward transfer for NLU/NLG datasets as the VL tasks are learned, which is rarely observed in CL. For the 160M parameter model, naive fine-tuning suffers large amounts of forgetting compared to mSGM. For some tasks, naive fine-tuning exhibits greater losses of NLU/NLG performance, unlike mSGM. In general, VL performance for mSGM rivals or exceeds naive fine-tuning across scales.
(a) Continual VL Performance
(b) Continual Linguistic Forgetting
Figure 5: Continual Learning LLaVA Tasks: Vision-language performance and Linguistic forgetting for varying model sizes. We evaluate mSGM + Rehearsal $(1\%)$ and LLaVA Naive-FT on the Continual Learning setup, with varying base LLMs: Pythia models from 160M to 1.4B scale. Task 2 is not associated with any VL dataset for evaluation, since it is a captioning and instruction following task.
C.3 Analysis of Forgetting across NLU and NLG tasks
@@
To understand the composition of linguistic forgetting on the LLaVA setting, we look at the NLU and NLG forgetting for 9 different LLMs with varying scales: Pythia scaling family of models from the 160M to the 2.8B parameter range, Phi2 3B, Vicuna1.5 7B and LLaMA2 7B. Figure  6 shows these results. We can observe a clear trend of higher forgetting for the NLG dataset (Lambada), compared to forgetting for NLU. Another trend we note is that forgetting typically reduces with higher model scales, for both NLG and NLU.
@@
Figure 6: Linguistic forgetting by NLU and NLG tasks: For the LLaVA setting, we look at different model scales and families, and show the degree of linguistic forgetting by both NLU and NLG tasks separately. Here, the NLU and NLG averages are computed as the simple mean.
C.4 Pre-trained LLMs
@@
C.5 Artifact Use
We ensure that any artifacts (such as datasets, software, models, code, or other supplementary materials) associated with our paper are used in a manner that aligns with their original purpose and the guidelines set forth by the creators. In particular, our artifacts are the models and code we build our experiments on top of, which we have listed above in Sections A.1 and C.4.

### v2

@@ C.2 Evaluation Dataset Splits
We use the slim [^3] https://github.com/TRI-ML/vlm-evaluation versions of all VQAv2, GQA, Text-VQA and RefCOCO datasets for evaluation, as provided within Prismatic-VLMs (Karamcheti et al., 2024). All slim versions of the evaluation sets contain $1024$ examples each, and we use the provided index splits for testing.
Appendix D Additional Experiments
D.1 Analysis of Forgetting across NLU and NLG tasks
@@
To understand the composition of linguistic forgetting on the LLaVA setting, we look at the NLU and NLG forgetting for 9 different LLMs with varying scales: Pythia family 160M to 2.8B, Phi2 3B, Vicuna1.5 7B and LLaMA2 7B. Figure  5 shows these results. Here, the NLU and NLG averages are computed as the simple mean. We can observe a clear trend of higher forgetting for the NLG dataset (Lambada), compared to forgetting for NLU. Another trend we note is that forgetting typically reduces with higher model scales, for both NLG and NLU.
@@
Figure 5: Linguistic forgetting by NLU and NLG tasks: For the LLaVA setting, we look at different model scales and families, and show the degree of linguistic forgetting by both NLU and NLG tasks separately.
D.2 Analysis of Linguistic Forgetting across multiple NLG tasks
In Figure  6, we analyze the degree of linguistic forgetting across multiple NLG datasets: WebQuestions (Berant et al., 2013), TriviaQA (Joshi et al., 2017) and Lambada (Paperno et al., 2016). We show for various methods using the LLaVA recipe at the 160M scale, that the degree of NLG forgetting is significant (between 40% and to 64%) for the Original LLaVA (Naive FT) across various NLG datasets, and is not specific to Lambada (Paperno et al., 2016). Additionally, we observe that our best method - Soft Targets outperforms other mitigation approaches by large margins in almost all cases, towards mitigating the degree of NLG forgetting. While LoRA and mSGM, appear to help mitigate forgetting, we hypothesize that the retention of lower probability logits in the soft targets approach is critical for retaining linguistic generation abilities.
Figure 6: Linguistic forgetting across various NLG tasks: For the LLaVA setting, we look at the degree of linguistic forgetting across multiple NLG datasets: WebQuestions (Berant et al., 2013), TriviaQA (Joshi et al., 2017) and Lambada (Paperno et al., 2016). In order to use a single metric across WebQuestions and TriviaQA (exact-match (EM)) and Lambada (accuracy), we use the “% Change” metric $\Delta^{\prime}=\frac{w_{\mathrm{(Base\;LLM)}}-w_{(\mathrm{MLLM\;Method})}}{w_{\mathrm{(Base\;LLM)}}}\times 100\%$. The NLG forgetting across all methods is reported at the 160M scale. The error bars show the degree of standard error in the NLG evaluations. Lower values are better. Negative $\Delta^{\prime}$ values indicate positive backward transfer.
D.3 Analyzing Task Order Sensitivity
We report the performance of the mSGM + Reh. and the Original LLaVA (Naive FT) for varying task orders in Tables  9 and 10. Naive FT, unlike mSGM + Reh., retains little to no VL performance by Task 5, with large NL forgetting.
Table 9: Task order 1: Task-wise Accuracies and Forgetting of Each Mitigation Method across VL and NL tasks
Model Task 2 (Instruct) Task 3 (OCR) Task 4 (Ref) Task 5 (VQA)
LLaVA Naive FT 0.10 3.72 0.10 17.17 0.12 12.71 0.14 26.61
mSGM + Reh. $(1\%)$ 0.10 -0.46 0.06 1.79 0.13 1.94 0.46 4.66
Table 10: Task order 2: Task-wise Accuracies and Forgetting of Each Mitigation Method across VL and NL tasks
Model Task 2 (VQA) Task 3 (OCR) Task 4 (Instruct) Task 5 (Ref)
LLaVA Naive FT 32.41 15.74 0.56 11.20 0.10 6.49 0.15 18.24
mSGM + Reh. $(1\%)$ 8.90 6.39 2.82 9.62 2.28 2.54 0.44 5.07
D.4 Task wise accuracies for Instruction-tuning Models
The individual task accuracy breakups for both NL as well as VL tasks are shown in Tables 11,  12. We observe as before instruction-tuned LLaMA 2 (7B) and Pythia (1.4B) models show minimal or even negative linguistic forgetting, indicating positive backward transfer. Conversely, base LLMs display greater variability, with base Pythia experiencing more forgetting than base LLaMA 2, possibly due to differences in their pre-training datasets affecting post-multimodal training linguistic performance. Instruction tuning significantly reduces NLG forgetting in Pythia (1.4B) and induces positive backward transfer from VL to NLG, with a similar trend observed in NLU tasks. This suggests that instruction-tuned LLMs exhibit little to no forgetting, likely due to the distributional similarity between text-only instruction tuning tasks and visual-instruction tuning, helping to mitigate catastrophic forgetting during multimodal training.
Table 11: Natural Language Task Accuracies: Instruction Tuning Analysis between Base LLMs vs LLaVA MLLMs. Raw accuracies on NL tasks are shown for our instruction‑tuning study. Instr. column is checked, if the underlying LLM is instruction‑tuned. Base LLMs are the underlying LLMs used to create the LLaVA MLLM.
Model Instr. WSC273 Winogrande ARC‑E ARC‑C Lambada
Base LLMs
Pythia (1.4B) 70.70 56.51 61.74 27.47 48.98
59.34 50.43 48.23 26.79 33.79
LLaMA2 (7B) 85.35 69.53 75.63 43.17 64.35
80.59 69.22 76.26 43.43 68.27
LLaVA MLLMs
Pythia (1.4B) 67.40 56.20 60.98 27.47 40.91
64.84 54.06 51.94 25.68 34.80
LLaMA2 (7B) 85.35 68.35 75.97 45.39 62.31
87.91 70.56 77.06 44.62 68.70
Table 12: Vision‑Language Task Accuracies: Instruction‑tuning Analysis. Raw accuracies on VL tasks are shown for our instruction‑tuning study. Instr. column is checked, if the underlying LLM is instruction‑tuned.
Model Instr. VQA TextVQA‑OCR TextVQA‑Pure GQA
Pythia (1.4B) 66.17 38.49 35.49 46.09
66.46 39.12 34.35 46.88
LLaMA2 (7B) 74.50 56.28 45.95 56.25
D.5 Task-wise Accuracies during Continual Training
In Table LABEL:tab:vl_acc_long, we report the raw accuracies during continual training. We observe that Original LLaVA (Naive FT) heavily biases on the first task that is learned i.e. VQAv2, and consequently performs strongly on the VQA evaluations (VQAv2 and GQA) as well as the TextVQA-Pure benchmark, which is also similarly a VQA task. However, it is unable to retain performance on OCR tasks after learning RefCOCO. Additionally, it completely loses its plasticity and fails to learn the RefCOCO task. As before heavily biases on the VQA tasks. The mSGM + Rehearsal($1\%$) approach on the other hand, learns and retains both OCR and VQA performances more effectively than all other approaches, while simultaneously also outperforming all other approaches on the final RefCOCO task - the most complex task in the LLaVA recipe for MLLMs to learn in our tests. This evidences the ability to effectively manage both plasticity and mitigating forgetting during the continual LLaVA training.
Table 13: Raw VL Accuracy during Continual LLaVA. We report the raw accuracies (%) on the five vision‐language tasks: VQAv2 (3), GQA (3), TextVQA-OCR (4), TextVQA Pure (4) and RefCOCO (5), evaluated after continual training up to each stage (2: Instruct, 3: VQA, 4: OCR, 5: Ref). The index of the evaluated vision-language tasks show which stage they are expected to evaluate for. As before, stage 1 refers to the LLM pre-training stage. Rows (“Trained on $\downarrow$”) indicates the last task stage seen during training and columns (“Evaluated $\rightarrow$”) denote the task under test. For “Instruct (2)”, the evaluation accuracies are reported where available. Each sub-table shows one mitigation method; stable accuracy across rows demonstrates effective knowledge retention, whereas drops on earlier tasks after later‐stage training reveal catastrophic forgetting. In reporting the task-averaged accuracies in our main results, we consider only the average of VL tasks it is trained on.
Trained on $\downarrow$ Evaluated $\rightarrow$
VQAv2 (3) GQA (3) TextVQA-OCR (4) TextVQA-Pure (4) RefCOCO (5)
Original LLaVA
Instruct (2)
VQA (3) 51.44 38.77 9.84 10.78 0.00
OCR (4) 45.13 30.47 9.86 11.96 0.00
Ref (5) 46.84 30.57 3.90 10.57 0.00
Soft Targets (ST)
Instruct (2) 0.45 0.00 1.55 1.15 0.00
VQA (3) 0.39 0.00 0.31 0.41 0.00
OCR (4) 35.64 26.07 8.03 5.00 0.00
Ref (5) 4.92 2.15 0.21 1.26 0.00
mSGM
Instruct (2) 0.10 0.00 0.13 0.00 0.00
VQA (3) 43.53 31.15 12.63 8.50 0.00
OCR (4) 36.11 26.66 6.90 7.57 0.00
Ref (5) 8.58 7.03 0.21 3.00 0.10
Rehearsal $(1\%)$
Instruct (2)
VQA (3) 45.52 32.23 6.59 9.36 0.00
OCR (4) 25.78 21.39 1.07 7.41 0.00
Ref (5) 23.49 14.75 1.42 6.87 2.25
mSGM + Rehearsal $(1\%)$
Instruct (2) 0.10 0.00 0.13 0.00 0.00
VQA (3) 42.85 29.98 13.41 8.35 0.00
OCR (4) 40.51 29.30 7.64 7.50 0.00
Ref (5) 39.66 28.81 13.89 7.99 4.30
Table 13: (continued) Raw VL Accuracy during Continual LLaVA
D.6 Pre-trained LLMs
@@
D.7 Artifact Use
We ensure that any artifacts (such as datasets, software, models, code, or other supplementary materials) associated with our paper are used in a manner that aligns with their original purpose and the guidelines set forth by the creators. In particular, our artifacts are the models and code we build our experiments on top of, which we have listed above in Sections D.6.